The times-interest-earned (TIE) ratio shows how well a firm can cover its interest payments with operating income.

Compare the income statements of Lost Pigeon Aviation and Purple Panda Importers and calculate the TIE ratio for each firm.

Lost Pigeon Aviation Income Statement For the Year Ended on December 31

(Millions of dollars)

Net Sales 1400
Variable costs 560
Fixed costs 490
Total Operating Costs $1,050.00
Operating Income (or EBIT) 350
Less interest 50
Earnings before Taxes (EBT) 300
Less taxes (40%) 120
Net Income 180
Times Interest Earned (TIE) ?

Purple Panda Importers Income Statement For the Year Ended on December 31

(Millions of dollars)

Net Sales $1,700.00
Variable costs 425
Fixed costs 765
Total Operating Costs $1,190.00
Operating Income (or EBIT) 510
Less interest 80
Earnings before Taxes (EBT) 430
Less taxes (40%) 172
Net Income 258
Times Interest Earned (TIE) ?

Complete the following statement, based on the calculations you have already made.

Describe the relationship between the TIE ratios of the two companies.

-Lost Pigeon Aviation has a greater TIE ratio than Purple Panda Importers.

-The companies have equal TIE ratios.

-Purple Panda Importers has a greater TIE ratio than Lost Pigeon Aviation.

Which company is in better position to cover its interest payments, and therefore exhibits lower risk, than the other?

-Purple Panda Importers is in a better position to cover its interest payment.

-Lost Pigeon Aviation is in a better position to cover its interest payment.

-Both companies are equally positioned to cover their interest payments.
Answer #1

TIE Ratio = EBIT/Interest

Lost Pigeon Aviation = 350/50 = 7 times

Purple Panda = 510/80 = 6.375

Higher the ratio, better it is

-Lost Pigeon Aviation has a greater TIE ratio than Purple Panda Importers

-Lost Pigeon Aviation is in a better position to cover its interest payment.